Problem 21.43

A circuit consists of a 12.0-V battery connected to three resistors (43? , 18? and 110? ) in series.

Part A

Find the current that flows through the battery.

Express your answer using two significant figures.

I =   mA  

SubmitMy AnswersGive Up

Part B

Find the potential difference across the 43? resistor.

Express your answer using two significant figures.

V1 =   V  

SubmitMy AnswersGive Up

Part C

Find the potential difference across the 18? resistor.

Express your answer using two significant figures.

V2 =   V  

SubmitMy AnswersGive Up

Part D

Find the potential difference across the 110? resistor.

Express your answer using two significant figures.

V3 =   V  

SubmitMy AnswersGive Up

Provide FeedbackContinue
Answer #1

A)
I=V/(R1+R2+R3)
=12/(43+18+110)
=0.07 A

B)
V1= I*R1
   = 0.07 * 43
= 3.01 V

C)
V2= I*R2
   = 0.07 * 18
= 1.26 V

D)
V3= I*R3
   = 0.07 * 110
= 7.7 V
Answer #2

A)current that flows through the battery I=V/Reff

Reff=R1+R2+R3

I=12/(43+18+110)=70 mA

B)potential difference across the 43 V43=I*43=70*10^-3*43=3.0V

C)potential difference across the 18 V18=I*18=70*10^-3*18=1.3V

D)potential difference across the 110 V110=I*110=70*10^-3*110=7.7V